# Cyber Defense Reinvented: Large Language Models as Threat Intelligence Copilots

### v1

@@ 5 Evaluation Results
Table 3: Evaluation results for ❷ Contextualization tasks, where "Hist."–historical threats, "0-Day"–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
@@
5.1 Threat Attribution
We begin by attributing the origin of cyber threats with three targets: threat actor, TTP, and exploitation path, as outlined in Table 1. Our results yield several research findings (RF):
RF1: Specialized Development is Necessary. As shown in Table 2, industry-leading LLMs, despite their broad applicability across diverse tasks, exhibit limited performance in CTI analysis. For instance, ChatGPT-4o has achieved only
SYSTEM You are a cybersecurity threat analyst specializing in cyber threat intelligence (CTI). Your task is to assess potential attack tactics, techniques, and procedures (TTPs), and provide an intelligence assessment following structured industry frameworks such as $\cdots$ USER <Report Title> "Emerging APT Group Exploiting Zero-Day Vulnerabilities in Cloud Infrastructure." <Report Excerpt> Recent intelligence reports indicate that an Advanced Persistent Threat (APT) group has been exploiting an unpatched zero-day vulnerability in cloud service environments. The group is using spear-phishing emails to gain initial access, deploying a custom malware variant named ‘CloudPhantom,’ and leveraging compromised API keys to move laterally $\cdots$ <Question> Identify the relevant TTPs used in this attack. Let’s think step-by-step, first, determine CTI tasks among $\cdots$ ASSISTANT <CTI Task><Analytical Target><Key Entities><Entity Relations><Retrieved Evidence and Source>$\cdots$<Final Answer>
Figure 4: Prompt example for TTP analysis. For both baselines and CyLens, we prompt them to perform step-by-step analysis following the NLP modules defined in Table 1. For tasks with prior dependencies (e.g., remediation), we instruct LLMs to first complete the prerequisite tasks before proceeding.
40.31% accuracy in threat actor attribution, significantly lower than CyLens-1B (79.52%) and CyLens-8B (83.74%) for the same task. This gap underscores the necessity of specialized development, as detailed in section 3 and section 4. Tailoring models specifically for CTI significantly reduces parameter requirements—CyLens-1B and 8B are far more lightweight compared to the 500B–1T parameters of GPT-4o.
RF2: Large model size is not always beneficial. Interestingly, scaling up model size does not always translate to improved CTI effectiveness. For instance, CyLens-8B achieves 93.74% recall in TTP attribution, outperforming CyLens-70B (88.26%), suggesting diminishing returns despite undergoing the same development process. This indicates that larger models do not alaways yield better performance in CTI tasks, emphasizing the importance of domain-specific optimizations over sheer model size.
RF3: Task-specific capability varies across baselines. As shown in Table 2, baseline models exhibit varying levels of effectiveness across different CTI tasks. For instance, GPT-o1 performs obviously worse in attributing threat actors compared to exploitation paths. A deeper case study reveals that baseline LLMs struggle to accurately interpret CTI-specific terminology related to threat actors. For example:
Case Study 1
@@ 5.1 Threat Attribution
On the other hand, GPT-o1 achieves $>$78% accuracy in analyzing exploitation paths, as this task primarily involves summarizing existing evidence instead of complex reasoning.
Case Study 2
In cases where specificity is less critical, a rough description—such as "JNDI lookup and remote code fetching"—sufficiently summarizes the Log4Shell exploitation path. While a more precise description would be: "The Log4j library contacts the attacker’s malicious LDAP server, which then responds with a reference to a remote Java class file containing malicious code," the broader summary still effectively conveys the core mechanism.
@@
RF4: CyLens Performs consistently across analyses. Unlike baseline models, CyLens maintains consistent performance across different analytical targets, highlighting the importance of incorporating comprehensive expertise during development. This balanced capability ensures that CyLens does not exhibit biased performance or knowledge gaps throughout the CTI lifecycle. The same conclusion is further reinforced by subsequent evaluations.
5.2 Threat Contextualization
We continue evaluating CyLens in contextualizing affected systems, attack infrastructure (e.g., domains and servers), and impact. Besides the similar findings as outlined in RF1-RF4, our evaluations derive complementary research findings:
RF5: Zero-day threat hunting is not always harder. Both Table 2 and 3 reveal that analyzing zero-day threats does not always result in worse performance compared to historical threats. For instance, when assessing the impact of cyber threats, LLMs such as CyLens-1B and SevenLLM sometimes achieve slightly better relevance in their generated analyses. We attribute this to the nature of specific CTI tasks: For impact analysis, the contextualized understanding of previously known threats allows LLMs to provide more insightful assessments, particularly when certain zero-day threats resemble threats they have already encountered. However, for tasks such as identifying affected systems, attack infrastructures, or threat actors, zero-day threats often introduce novel and previously unseen products or malware, making analysis more challenging compared to historical threats.
RF6: Larger LLMs tend to be wordy. Through detailed case study, we observe that larger LLMs tend to generate verbose responses, leading to lower text similarity as measured by BERT score. For example:
Case Study 3
CyLens-8B summarizes an impact analysis concisely: "The March 2022 cyberattack by the Lazarus Group resulted in the theft of \$620 million from the Ronin Network, significantly impacting the cryptocurrency market and…" This closely aligns with the reference [101], maintaining factual accuracy while remaining succinct. In contrast, CyLens-70B generates a more verbose response: "In March 2022, the notorious Lazarus Group orchestrated a sophisticated cyber heist, infiltrating the Ronin Network’s defenses to abscond with a staggering \$620 million in cryptocurrency. This audacious act not only sent shockwaves through the digital currency markets, causing a temporary plunge in asset values, but also undermined investor confidence…" The latter output is overly wordy and includes sentiment-heavy language (e.g., "notorious," "sophisticated," "staggering," "audacious"), making it less aligned with the original reference.
@@ 5.2 Threat Contextualization
RF7: CTI agents should align with real-world practice. Despite the strong performance of SOTA agents (SevenLLM and Lily-Cyber) on certain tasks (e.g., analyzing TTPs), they struggle with others, such as attack infrastructure analysis. We attribute this to a misalignment between their training focus and CTI applications. For instance, SevenLLM is trained with a broad emphasis on NLP tasks (e.g., NER), while Lily-Cyber is developed using hand-crafted cybersecurity and hacking-related data pairs [84]. However, these training paradigms often conflate TTPs and threat actors, rather than treating them as distinct analytical components. This lack of granularity can hinder precise CTI reasoning, particularly for tasks that require structured analysis of attack properties.
Although CTI lacks a universally standardized taxonomy, structuring tasks in alignment with real-world CTI databases (e.g., MITRE-CVE and NVD) provides a more practical and reliable foundation for cyber threat analysis. This highlights the necessity of LLMs specifically designed for CTI, rather than relying on models with indirect exposure to cybersecurity knowledge or misaligned customization that does not fully reflect real-world CTI practices.
Table 4: Evaluating ❺ Prioritization performance using CVSS 3.x metrics [60]: AV–Attack Vector, AC–Attack Complexity, PR–Privileges Required, UI–User Interaction, S–Scope, C–Confidentiality Impact, I–Integrity Impact, A–Availability Impact, Base–CVSS Base Severity. We report the accuracy of correctly categorizing each metric (e.g., Attack Vector: Network). The boldface and underline highlight the best and second-best results, respectively.
@@
5.3 Detection and Correlation
We now evaluate CyLens and baselines on ❸ detection and ❹ correlation tasks. In detection, the objective is to evaluate how effectively LLMs can identify relevant threat identifiers (CVEs and CWEs) from a given threat report. This step is crucial for enabling subsequent CTI activities, such as threat correlation and remediation. In correlation, we extend the evaluation by examining how LLMs associate other threats (CVEs) using different strategies: (i) Direct correlation–directly correlate additional CVEs. (ii) CVEs as intermediates–first detect CVEs from threat reports (from ❸) and then use them for correlation. (iii) CWEs as intermediates–first identifying CWEs (i.e., weakness categories, from ❸) and using them for correlations. The latter two strategies leverage intermediate evidence (CVEs or CWEs identified in the detection step) to uncover the threat network.
Our evaluations yield the following new research findings:
Figure 5: Evaluation results on ❸ Detection tasks using historical threats. The presentation order of LLM names (top-down) follows the same sequence as in Tables 2 (top-down).
RF8: CyLens demonstrates strong detection relevance. Figure 5-(a)(b) shows the hit ratio for the top 10 detected threat identifiers (CVEs, CWEs). Identifying CVEs is more challenging due to their vast number ($>$200k) compared to CWEs ($\sim$900). However, we also observe that CyLens  is able to detect more relevant CVEs such as those that belong to the same category or affect the same products. For example:
Case Study 4
Given the threat report: "…Apache HTTP Server allows remote attackers to execute arbitrary code. The issue arises from improper handling of user-supplied input in HTTP request headers…" corresponding to CVE-2021-42013. LLMs present different detections: GPT-o1 and DeepSeek-R1 incorrectly output CVE-2020-0609 and CVE-2022-2296, both of which are unrelated to Remote Code Execution or Apache. CyLens-70B gives CVE-2023-25690, a highly relevant HTTP Request Smuggling attack on the Apache Server.
From the above case study, although CyLens-70B does not match the exact CVE identifier, it effectively aligns with the evidence provided in the threat report and demonstrate a strong threat detection capability. This effectiveness is further supported by the results shown in Figure 5-(c)(d).
RF9: CyLens demonstrates robust threat correlation. Table 5 presents the hit ratio for the top 10 correlated threats (CVEs) as reported in MITRE-CVE [14]. We evaluate three correlation strategies: (i) Direct correlation, (ii) CVEs as intermediates, and (iii) CWEs as intermediates, as described above. We have corresponding observations:
Table 5: Evaluation results (HIT@10) on ❹ Correlation task using historical threats. The order of LLM names (left-to-right) follows the same sequence as in Tables 2 (top-down).
@@ 5.3 Detection and Correlation
      While baseline models largely struggle to correlate CVEs, CyLens demonstrates significant effectiveness (e.g., 77.5% Hit@10 with CyLens-70B).
@@
      Incorporating CVEs as intermediate evidence improves correlation performance. However, baseline models tend to propagate erroneous CVEs from the prior detection step, which misguides subsequent threat correlations.
@@
      Incorporating CWEs degrades correlation performance due to their inherent ambiguity. Since a CWE can broadly relate to many CVEs, it introduces ambiguity, impeding accurate threat correlation. Despite this challenge, CyLens remains robust (e.g., 71.6% Hit@10 with CyLens-70B) in incorporating ambiguous intermediate evidence while maintaining strong correlation effectiveness.
5.4 Threat Prioritization
@@
Real-world cyber threats often co-occur across cyber events [77, 73], necessitating that CTI teams distinguish threat severity and prioritize subsequent patches and mitigations. Our evaluation focuses on two prioritization measurements: severity level and analysis of exploitation dynamics.
@@
Severity Level. Threat severity can be assessed from threat reports even when CVE identifiers are not disclosed or when zero-day threats have not yet been assigned CVEs. We evaluate severity using 8 metrics defined in CVSS 3.x [60], which collectively capture a threat’s exploitability and impact. Each severity metric is treated as a classification task. For example, the Attack Vector metric has four predefined categories: Network, Adjacent Network, Local, and Physical. In cases where a threat report lacks sufficient evidence for CVSS metrics, we introduce an additional category, "N/A", for each metric.
@@
Furthermore, we classify the base severity level from four categories: Low, Medium, High, and Critical. This classification aggregates the ratings from 8 CVSS metrics, providing an overall severity assessment aligned with CVSS standards.
@@
RF10: CyLens maintains stable effectiveness. Table 4 shows the accuracy of all models. While baselines vary widely across CVSS metrics (e.g., GPT-o1 scores 92.91% on Attack Complexity but only 20.43% on Integrity Impact), CyLens remains consistently strong. Notably, CyLens-8B achieves $>$90% accuracy across 8 metrics for historical threats, demonstrating its reliability in handling different severity factors.
@@
This stability comes from CyLens ’s ability to adapt to different evidence types. Each CVSS metric relies on specific context—Privileges Required (PR) depends on access control policies, while Attack Complexity (AC) is influenced by real-world constraints and attacker skill level. By effectively integrating diverse evidence sources, CyLens reduces inconsistencies in severity analyses, which aligns with RF4.
Additionally, we find that zero-day threats are not always harder to analyze, aligning with RF5. Some evidence, like some user actions in exploitations, is already known from past threats, facilitating assessment on newly emerging threats.
We also observe that larger LLMs do not always perform better, consistent with RF6. While bigger models tend to provide more detailed reasoning, this can sometimes lead to inaccurate analysis. For example, CyLens-70B does not always outperform CyLens-8B across all severity metrics.
@@
Exploitation Dynamics. Besides severity analysis using CVSS, we also evaluate CyLens’s effectiveness in analyzing dynamic exploitation behaviors. Specifically, we use the EPSS score [29] and conduct two types of analysis: (1) pinpointing historical exploitation at an earlier timestamp and (2) forecasting future exploitation at a later timestamp.
@@
Figure 6: With $\mathcal{T}_{0}$ as the “current” timestamp (varying among threats), we compute RMSE for historical EPSS scores over the past three months (interpolation) and future EPSS scores for the next three months (prediction) w.r.t. $\mathcal{T}_{0}$.
@@
We compare CyLens-1B/8B against four baseline models and compute RMSE between the interpolated (historical) or predicted (future) EPSS scores and the actual scores (scaled from 0% to 100%). To ensure accurate evaluation, we match each EPSS score to the closest available recorded score. For example, if an EPSS score is not reported exactly on 10/01/2023, we use the closest available date (e.g., 09/25/2023) from EPSS reports [14] as the ground truth.
RF11: LLMs vary in interpolation and prediction. As shown in Figure 6, LLMs perform slightly better in interpolating historical EPSS scores (lower RMSE) than in predicting future scores. However, we also observe that local LLMs, such as SevenLLM, perform obviously worse in both interpolation and prediction due to their static inherent knowledge, which lacks adaptability to evolving cyber threats. On the contrary, CyLens, despites its local deployment, performs similar to API-based LLMs (e.g., Gemini), which demonstrates the necessity of equipping LLMs with domain expertise. This enables LLMs to adapt to the dynamic nature of CTI tasks.
RF12: Contextual knowledge is essential. Beyond Figure 6, we observe that some cyber threats undergo dramatic shifts in EPSS scores due to emerging exploitation activities.
Case Study 5
For instance, CVE-2024-27956 saw a sharp increase in EPSS on 02/15/2025, jumping from 0.31% to 55.37% [19]. No LLM was able to accurately predict this trend without contextual knowledge of real-world events.
This implies the importance of not only fitting to EPSS trends but also incorporating real-world evidence in real time to enhance CTI adaptability.
5.5 Remediation
Table 6: Evaluation results on ❻ Remediation. We evaluate whether LLMs can correctly generate: (i) patching or mitigation tools – measured by accuracy, (ii) code patches – measured by syntactic similarity, (iii) patching or mitigation methods – measured by BERT Score, and (iv) reference advisories – measured by accuracy.

### v2

@@ 5. Evaluation Results: Broad CTI
Table 5. Evaluation results for ❷ Contextualization tasks, where ”Hist.”–historical threats, ”0-Day”–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
@@
5.1. Threat Attribution
We begin by attributing the origin of cyber threats with three targets: threat actor, TTP, and campaign, as outlined in Table 2. Our results yield several research findings (RF):
RF1: Specialized development is necessary. As shown in Table 3, industry-leading LLMs, despite their broad applicability across diverse tasks, exhibit limited performance in CTI analysis. For instance, ChatGPT-4o has achieved only 40.31% accuracy in threat actor attribution, significantly lower than CyLens-1B (79.52%) and CyLens-8B (83.74%) for the same task. This gap underscores the necessity of specialized development, as detailed in section 3. Tailoring models specifically for CTI significantly reduces parameter requirements—CyLens-1B and 8B are far more lightweight compared to the 500B–1T parameters of GPT-4o.
RF2: Large model size is not always beneficial. Interestingly, scaling up model size does not always translate to improved CTI effectiveness. For instance, CyLens-8B achieves 93.74% recall in TTP attribution, outperforming CyLens-70B (88.26%), suggesting diminishing benefits. This indicates the importance of domain-specific optimizations instead of sheer model sizes.
RF3: Task-specific capability varies across baselines. As shown in Table 3, baseline models exhibit varying levels of effectiveness across different CTI tasks. For instance, GPT-o1 performs obviously worse in attributing threat actors compared to campaign. A deeper case study reveals that baseline LLMs struggle to accurately interpret CTI-specific terminology related to threat actors. For example:
Case Study 1.
@@ 5.1. Threat Attribution
On the other hand, GPT-o1 achieves $>$78% accuracy in analyzing attack campaign, as this task primarily involves summarizing existing evidence instead of complex reasoning.
Case Study 2.
In cases where specificity is less critical, a rough description—such as ”JNDI lookup and remote code fetching”—sufficiently summarizes the campaign of Log4Shell. While a more precise description would be: ”The Log4j library contacts the attacker’s malicious LDAP server, which then responds with a reference to a remote Java class file containing malicious code,” the broader summary still effectively conveys the core mechanism.
@@
5.2. Threat Contextualization
We continue evaluating CyLens in contextualizing affected systems, attack infrastructure (e.g., domains and servers), and impact.Our evaluations derive complementary research findings:
RF4: Larger LLMs tend to be wordy. From Table 5, we observe that larger LLMs tend to achieve lower text similarity scores as measured by BERTScore. For example, CyLens-70B performs worse than CyLens-8B in generating threat impact descriptions on historical threats (87.33 vs. 90.03). Through detailed case studies, we find that larger LLMs are more likely to produce verbose responses:
Case Study 3.
CyLens-8B summarizes an impact analysis concisely: ”The March 2022 cyberattack by the Lazarus Group resulted in the theft of \$620 million from the Ronin Network, significantly impacting the cryptocurrency market and…” This closely aligns with the reference (Wikipedia contributors, 2025b), maintaining factual accuracy while remaining succinct. In contrast, CyLens-70B generates a more verbose response: ”In March 2022, the notorious Lazarus Group orchestrated a sophisticated cyber heist, infiltrating the Ronin Network’s defenses to abscond with a staggering \$620 million in cryptocurrency. This audacious act not only sent shockwaves through the digital currency markets, causing a temporary plunge in asset values, but also undermined investor confidence…” The latter output is overly wordy and includes sentiment-heavy language (e.g., ”notorious,” ”sophisticated,” ”staggering,” ”audacious”), making it less aligned with the original reference.
@@ 5.2. Threat Contextualization
RF5: Zero-day threat hunting is not always harder. Both Table 3 and 5 reveal that analyzing zero-day threats does not always result in worse performance compared to historical threats. For instance, when assessing the impact of cyber threats, LLMs such as CyLens-1B and SevenLLM sometimes achieve slightly better relevance in their generated analyses. We attribute this to the nature of specific CTI tasks: For impact analysis, the contextualized understanding of previously known threats allows LLMs to provide more insightful assessments, particularly when certain zero-day threats resemble threats they have already encountered. However, for tasks such as identifying affected systems, attack infrastructures, or threat actors, zero-day threats often introduce novel and previously unseen products or malware, making analysis more challenging compared to historical threats.
RF6: CTI agents should align with real-world practice. Despite the strong performance of SOTA agents (SevenLLM and Lily-Cyber) on certain tasks (e.g., analyzing TTPs), they struggle with others, such as attack infrastructure analysis. We attribute this to a misalignment between their training focus and CTI applications. For instance, SevenLLM is trained with a broad emphasis on NLP tasks (e.g., NER), while Lily-Cyber is developed using hand-crafted cybersecurity and hacking-related data pairs (Segolily Labs, 2025). However, these training paradigms often conflate TTPs and threat actors, rather than treating them as distinct analytical components. This lack of granularity can hinder precise CTI reasoning, particularly for tasks that require structured analysis of attack properties.
Table 6. Evaluating ❺ Prioritization performance using CVSS 3.x metrics (National Institute of Standards and Technology (2025), NIST): AV–Attack Vector, AC–Attack Complexity, PR–Privileges Required, UI–User Interaction, S–Scope, C–Confidentiality Impact, I–Integrity Impact, A–Availability Impact, Base–CVSS Base Severity. We report the accuracy of correctly categorizing each metric (e.g., Attack Vector: Network). The boldface and underline highlight the best and second-best results, respectively.
@@
5.3. Threat Correlation
We now evaluate CyLens on ❸ correlation tasks. The objective is to measure how effectively CyLens can identify relevant threat identifiers (CVEs and CWEs) from a given threat report. Moreover, we extend the correlation by identifying additional threats (CVEs) that are not directly from threat reports but are related or concurrent, i.e., the Expansive Correlation. We use different strategies: (i) Directly conduct expansive correlation. (ii) First find CVEs from threat reports and then expand to additional CVEs. (iii) First find CWEs from threat reports and then expand to additional CVEs. Our evaluations yield the following findings:
Figure 5. Evaluation results on ❸ Correlation using historical threats (we do not evaluate on 0-day threats as they have incomplete correlation reports). The presentation order of LLM names (top-down) follows the same presentation sequence as in Tables 3 (top-down).
RF7: CyLens demonstrates strong correlation relevance. Figure 5-(a)(b) shows the hit ratio for the top 10 correlated threat identifiers (CVEs, CWEs). Correlating CVEs is more challenging due to their vast number ($>$200k) compared to CWEs ($\sim$900). However, we also observe that CyLens  is able to correlate more relevant CVEs such as those that belong to the same category or affect the same products. For example:
Case Study 4.
Given the threat report: ”…Apache HTTP Server allows remote attackers to execute arbitrary code. The issue arises from improper handling of user-supplied input in HTTP request…” is related to CVE-2021-42013. LLMs present different outputs: GPT-o1 and DeepSeek-R1 incorrectly output CVE-2020-0609 and CVE-2022-2296, both of which are unrelated to Remote Code Execution or Apache. CyLens-70B gives CVE-2023-25690, a highly relevant HTTP Request Smuggling attack on the Apache Server.
From the above case study, although CyLens-70B does not match the exact CVE identifier, it effectively aligns with the evidence provided in the threat report and demonstrate a strong threat correlation capability. This effectiveness is further supported by the results shown in Figure 5-(c)(d), where CyLens presents stronger correlation relevance w.r.t. the correct products and threat categories.
RF8: CyLens demonstrates robust expansive correlation. Table 7 presents the hit ratio for the top 10 expanded threats (CVEs) as reported in MITRE-CVE (Corporation, nd). We evaluate three correlation strategies: (i) Direct expansion, (ii) CVEs as intermediates, and (iii) CWEs as intermediates. We have corresponding observations:
Table 7. Evaluation results (HIT@10) on expansive correlation task using historical threats (we do not evaluate on 0-day threats as they have incomplete reports of concurrent CVEs). The order of LLM names (left-to-right) follows the same presentation sequence as in Tables 3 (top-down).
@@ 5.3. Threat Correlation
      While baseline models largely struggle to expansively correlate CVEs, CyLens demonstrates significant effectiveness (e.g., 77.5% HIT@10 with CyLens-70B).
@@
      Incorporating CVEs as intermediate evidence improves expansive correlation. However, baseline models tend to propagate erroneous CVEs from the prior detection step, which misguides subsequent threat correlations.
@@
      Incorporating CWEs degrades expansive correlation due to their inherent ambiguity. Since a CWE can broadly relate to many CVEs, it introduces ambiguity, impeding accurate correlation.
5.4. Threat Prioritization
@@
Real-world cyber threats often co-occur across cyber events (Rahman and Williams, 2022; Peng et al., 2020), necessitating that CTI teams to distinguish threat severity and prioritize subsequent patches and mitigations. Our evaluation focuses on two prioritization measurements: severity level and analysis of exploitation dynamics.
@@
Severity Level. Threat severity can be assessed from threat reports even when CVE identifiers are not disclosed or when zero-day threats have not yet been assigned CVEs. We evaluate severity using 8 metrics defined in CVSS 3.x (National Institute of Standards and Technology (2025), NIST), which collectively capture a threat’s exploitability and impact. Each severity metric is treated as a classification task. For example, the Attack Vector metric has four predefined categories: Network, Adjacent Network, Local, and Physical. If a threat report lacks sufficient evidence for CVSS metrics, we introduce an additional label, ”N/A”, for each metric.
@@
Furthermore, we classify the base severity level from four categories: Low, Medium, High, and Critical. This classification aggregates the ratings from eight CVSS metrics, providing an overall severity assessment aligned with CVSS standards.
@@
RF9: CyLens maintains stable effectiveness. Table 6 shows the accuracy of all models. While baselines vary widely across CVSS metrics (e.g., GPT-o1 scores 92.91% on Attack Complexity but only 20.43% on Integrity Impact), CyLens remains consistently strong. Notably, CyLens-8B achieves $>$85% accuracy across 8 metrics, demonstrating its reliability across different severity factors.
@@
This stability comes from CyLens ’s adaptability across different evidence types. Each CVSS metric relies on specific evidence, for example, Privileges Required (PR) depends on access control policies, while Attack Complexity (AC) is influenced by system constraints and attacker’s skill. CyLens reduces inconsistencies in severity analyses by effectively integrating diverse evidence sources.
@@
Exploitation Dynamics. Besides severity analysis using CVSS, we also evaluate CyLens’s effectiveness in analyzing dynamic exploitation behaviors. Specifically, we use the EPSS score (FIRST.org, 2025) and conduct two types of analysis: (1) pinpointing historical exploitation at an earlier timestamp and (2) forecasting future exploitation at a later timestamp. Note that we exclude zero-day threats from the exploitation analysis due to the incompleteness of EPSS data.
@@
Figure 6. With $\mathcal{T}_{0}$ as the “current” timestamp (varying among threats), we compute RMSE for historical EPSS scores over the past three months (interpolation) and future EPSS scores for the next three months (prediction) w.r.t. $\mathcal{T}_{0}$.
@@
We compare CyLens-1B/8B against four baseline models and compute RMSE between the interpolated (historical) or predicted (future) EPSS scores and the actual scores (scaled from 0% to 100%). To ensure accurate evaluation, we match each EPSS score to the closest available recorded score. For example, if an EPSS score is not reported exactly on 10/01/2023, we use the closest available date (e.g., 09/25/2023) from EPSS reports (Corporation, nd) as the ground truth.
RF10: LLMs vary in interpolation and prediction. As shown in Figure 6, LLMs perform slightly better in interpolating historical EPSS scores (lower RMSE) than in predicting future scores. However, we also observe that local LLMs, such as SevenLLM, perform obviously worse in both interpolation and prediction due to their static inherent knowledge, lacking adaptability to the emerging exploitation. On the contrary, CyLens performs similar to API-based LLMs (e.g., Gemini), which demonstrates the necessity of equipping LLMs with domain expertise. This enables LLMs to adapt to the dynamic nature of CTI tasks.
RF11: Contextual awareness is required. We also observe that some cyber threats undergo dramatic shifts in EPSS scores due to emerging exploitation activities. This implies the importance of not only fitting to EPSS trends but also incorporating real-world events to enhance CTI capability. For instance:
Case Study 5.
CVE-2024-27956 presents a sharp increase in EPSS scores on 02/15/2025, jumping from 0.31% to 55.37% (CVE Details, 2024). No LLM was able to accurately predict this trend without contextual knowledge of real-world events.
Figure 7. Evaluation results by different groups of threats: (a)–(d) correspond to threats targeting the same product, (e)–(h) to those affecting the same vendor, and (i)–(l) to threats within the same category. We further divide these threats into historical and zero-day groups, with dataset statistics provided in Table 4. The evaluation metrics used are accuracy for both threat actor attribution and third-party advisory suggestions. Abbrevation: ”Bd CTI” – CyLens fine-tuned by broad CTI tasks (same setting as section 5); ”Sp CTI” – CyLens fine-tuned by specialized CTI tasks. Additional evaluation results are shown in Figure 8.
5.5. Remediation
Table 8. Evaluation results on ❻ Remediation. We evaluate whether LLMs can correctly generate: (i) patching or mitigation tools – measured by accuracy, (ii) code patches – measured by syntactic similarity, (iii) patching or mitigation methods – measured by BERT Score, and (iv) reference advisories – measured by accuracy.